OpenStudy (anonymous):

Express the complex number in trigonometric form. -2i

OpenStudy (anonymous):

easier than it looks there is no real part, only complex part, so the angle is \(-\frac{\pi}{2}\)

OpenStudy (anonymous):

oops angle is \(-\frac{\pi}{2}\) or if you prefer \(\frac{3\pi}{2}\)

OpenStudy (anonymous):

These are my answer choices 2(cos 180° + i sin 180°) 2(cos 90° + i sin 90°) 2(cos 270° + i sin 270°) 2(cos 0° + i sin 0°)

OpenStudy (anonymous):

and of course \(r=|-2i|=2\) you get \[-2i=2\left(\cos(\frac{3\pi}{2})+i\sin(\frac{3\pi}{2})\right)\]

OpenStudy (anonymous):

you are using degrees for this??? ok

OpenStudy (anonymous):

yeah degrees. my online teacher makes no sense

OpenStudy (anonymous):

convert \(\frac{3\pi}{2}\) to degrees

OpenStudy (anonymous):

270?

OpenStudy (anonymous):

so this would be the answer? 2(cos 270° + i sin 270°)